### v1

# Guiding diffusion models to reconstruct flow fields from sparse data

Abstract
The reconstruction of unsteady flow fields from limited measurements is a challenging and crucial task for many engineering applications. Machine learning models are gaining popularity in solving this problem due to their ability to learn complex patterns from data and generalize across diverse conditions. Among these, diffusion models have emerged as particularly powerful in generative tasks, producing high-quality samples by iteratively refining noisy inputs. In contrast to other methods, these generative models are capable of reconstructing the smallest scales of the fluid spectrum. In this work, we introduce a novel sampling method for diffusion models that enables the reconstruction of high-fidelity samples by guiding the reverse process using the available sparse data. Moreover, we enhance the reconstructions with available physics knowledge using a conflict-free update method during training. To evaluate the effectiveness of our method, we conduct experiments on 2 and 3-dimensional turbulent flow data. Our method consistently outperforms other diffusion-based methods in predicting the fluid’s structure and in pixel-wise accuracy. This study underscores the remarkable potential of diffusion models in reconstructing flow field data, paving the way for their application in Computational Fluid Dynamics research.
I Introduction
Fluid dynamics plays an essential role in many modern engineering applications, including vehicle and turbine design Pfeiffer (2016); Strom, Brunton, and Polagye (2017); Maurya, Maurya, and Luby (2018), active flow control Manohar et al. (2018); Brunton and Noack (2015); Sipp and Schmid (2016), cardiac blood flow modelling Yakhot, Anor, and Karniadakis (2007); Sankaran et al. (2012), ship wake detection Graziano, D’Errico, and Rufino (2016) and climate predictions Kalnay (2003). Numerical simulations and experimental studies are key methods of investigating complex flow phenomena. However, both approaches have inherent limitations. Although advancements in computational fluid dynamics (CFD) and hardware capabilities now enable high-fidelity simulations PI et al. (2013), accurate simulation of large-scale industrial problems still requires significant computational resources Carlberg et al. (2019). Thus, reconstructing fine flow structures from sparse data is still a critical topic for fluid simulations. Meanwhile, experimental studies are often limited by sensor numbers in the flow field, and evaluating complex flow behaviour through sparse signals from sensors is a huge challenge for experimental fluid dynamics Qin et al. (2022); Raffel et al. (2018).
Sparse reconstruction is a technique employed to recover or approximate the full-scale features of a system using only a limited subset of data, such as selected measurements, pixels, or observations. It has been pivotal in aerodynamic applications, as seen in early work of Bui-Thanh et al. Bui-Thanh, Damodaran, and Willcox (2004), where they utilized Proper Orthogonal Decomposition (POD) to resolve complex flow fields and aerodynamic forces efficiently. Modifications of this method like Gappy-POD Venturi and Karniadakis (2004); Raben, Charonko, and Vlachos (2012) or other linear methods such as Dynamic Mode Decomposition (DMD) Schmid (2010); Tu (2013) have been used to extract the coherent structures and correlation features of a fluid flow, which can later be used to reconstruct its high-resolution representation. A different approach was used in the work of Callaham et al. Callaham, Maeda, and Brunton (2018), where sparse representation was leveraged to reconstruct target states from a dictionary of precomputed examples, showing better robustness than the previously introduced methods when encountering noisy data. Loiseau et al. Loiseau, Noack, and Brunton (2018) presented a sparse reduced-order modeling framework that combined feature extraction, sparse nonlinear dynamics identification, and modal decomposition. They used it to reconstruct full-state fluid dynamics from typical experimental data such as time-resolved sensor data and optionally also non-time-resolved Particle Image Velocimetry (PIV) snapshots. The article from Zaki et al. Zaki and Wang (2021) proposed a hybrid approach bridging physics-based modelling and data-driven optimization, where an adjoint-variational method is introduced to reconstruct turbulent flow.
Meanwhile, the success of machine learning methods has motivated several studies addressing the challenges in data reconstruction with established machine learning pipelines. Deng et al. Deng et al. (2019) tried reconstructing flows using discrete point measurements and non-time-resolved PIV measurements based on long short-term memory (LSTM) networks. In the work of Erichson et al. Erichson et al. (2020), Shallow Neural Networks (SNN) were explored to reconstruct a flow field from sparse sensor measurements. This same architecture was leveraged to reconstruct a flow over a stalled aerofoil in the work of Carter et al. Carter et al. (2021), combining it with linear state estimation methods based on classical compressed sensing and extended POD methodologies. In this same field of aerodynamics, Zhong et al. Zhong et al. (2023) used a Convolutional Neural Network (CNN) combined with the multi-layer perceptron (MLP) to reconstruct the vorticity field surrounding the airfoil. Dubois et al. Dubois et al. (2022) compared several machine learning techniques for flow data reconstructions with limited measurements, demonstrating the potential of deep learning methods in the rapid estimation of complex flows.
Solving the sparse reconstruction problem can also be seen as a super-resolution task, where our set of sparse sensor measurements compose a low-resolution sample, and we aim to reconstruct its high-resolution counterpart Fukami, Fukagata, and Taira (2023). Sun et al. Sun and Wang (2020) proposed a physics-constrained Bayesian neural network (PC-BNN) to reconstruct a flow field, modelling their problem as a super-resolution task. Fukami et al. Fukami et al. (2021) took this same approach by computing a Voronoi tessellation of the sparse information and then using it as an input for a CNN, an established architecture for image super-resolution Dong et al. (2015). Another example is the work of Güemes et al. Güemes, Sanmiguel Vila, and Discetti (2022), where they presented a framework called RaSeedGAN that interpolated the scattered data onto a regular grid and afterwards upsampled it using a Generative Adversarial Network (GAN).
These methods have demonstrated remarkable performance in reconstructing fluid flows within the specific scenarios and parameter ranges they were designed to address. However, their effectiveness often diminishes when applied to cases that differ from their training data, such as flows with significantly different Reynolds numbers or unseen flow configurations. As a result, there is a need for approaches that are accurate within their domain and robust across a wide range of flow conditions and problem settings.
In recent years, diffusion models Sohl-Dickstein et al. (2015); Song et al. (2020); Ho, Jain, and Abbeel (2020) have shown superior performance in tasks like image generation Dhariwal and Nichol (2021), image editing Meng et al. (2021); Avrahami, Lischinski, and Fried (2022), image inpainting Lugmayr et al. (2022); Chung, Sim, and Ye (2022), and image colorization Saharia et al. (2022a); Yue, Wang, and Loy (2024). Following this success, there has been a rising interest in solving image super-resolution tasks with these models to improve in reconstruction accuracy. A direct approach to achieve this goal is to give the low-resolution image as an input to the model, and train it to map the high-resolution counterpart Saharia et al. (2022b); Rombach et al. (2022). Other approaches utilize a pre-trained diffusion model and only modify the sampling methods Chung, Sim, and Ye (2022); Choi et al. (2021); Yue and Loy (2024); Wang et al. (2024). More recently, diffusion models have also been used to predict physical flows. Liu et al. Liu and Thuerey (2024) explored their use as uncertainty-aware surrogate models for turbulence simulations, demonstrating their ability to capture full solution distributions and outperform traditional probabilistic approaches. Moreover, Shu et al. Shu, Li, and Farimani (2023) presented a super-resolution method for reconstructing high-fidelity flow fields using this same architecture. Diffusion models have also been used to predict nonlinear fluid fields from discrete simulation data Yang and Sommer (2023), turbulent rotating flows Li et al. (2023), and supersonic flows Abaidi and Adams (2025).
With their generative capabilities and the proficiency to accurately learn complex distributions, diffusion models emerged as a promising candidate to enhance the existing reconstruction methods found in the literature. They are especially interesting in combination with physical priors, such as conservation laws.
In terms of model architectures, early approaches primarily relied on CNNs applied to cartesian grids Thuerey et al. (2020). Later, neural operators emerged, most notably in the form of Fourier Neural Operators (FNO) Li et al. (2021), and Graph Neural Networks (GNNs) Battaglia et al. (2016); Lino et al. (2022). More recently, spatial transformer architectures were employed as neural operators, e.g., via OFormer Li, Meidani, and Farimani (2022), Universal Physics Transformers (UPT) Alkin et al. (2025) and Transolver Wu et al. (2024), among others. Transformers have experienced particular interest due to their ability to capture global and multi-scale interactions. Their main drawback is the computational $N^{2}$ scaling of their inherent attention mechanism, where N is the number of grid nodes. Thus, vanilla attention can be prohibitive for large 2D or 3D problems. The P3D model Holzschuh et al. (2025), which we employ in our experiments, employs shifted-window attention Liu et al. (2021) to improve on this behavior. Even with architectural advances, the computational cost of handling high-resolution 3D fields has restricted applications to low-resolution 3D settings with limited accuracy. This includes deterministic models Zhou et al. (2022); Maurya (2025), GANs Subramaniam et al. (2020); Kim et al. (2021), and diffusion approaches Bi et al. (2025); Huang et al. (2024). In this context, our work advances diffusion-based super-resolution methods to 3D fields with high resolutions.
In the following, we build on state-of-the-art diffusion modeling techniques to develop a novel algorithm for sparse data reconstruction tasks of fluids that outperforms existing methods. Specifically, we first introduce a masking approach that improves the denoising step of diffusion models. This masking is combined with a modified, conflict-free update step to ensure that disagreements in the learning feedback from denoising and physics constraints do not hinder learning. We will demonstrate the gains in distributional accuracy and obeying physics constraints for a series of challenging fluid sparse data reconstruction tasks with turbulent flows.
II Methodology
In this section, we will briefly introduce diffusion models to establish the theoretical framework required to present our new sampling method. Furthermore, we discuss the approaches used to train the underlying deep neural networks to ensure physical consistency.
II.1 Diffusion models
In this work, we aim to reconstruct a high-fidelity sample $\mathbf{x_{0}}\in\mathbb{R}^{n\times n}$ from a subset of sparse true values $\mathbf{x_{s}}$ with denoising diffusion probabilistic models (DDPMs) Sohl-Dickstein et al. (2015); Ho, Jain, and Abbeel (2020). DDPMs are generative models that learn to gradually transform samples from a simple known distribution $p(\mathbf{x}_{T})$ into samples from the target distribution $p(\mathbf{x}_{0})$ that we want to model. Typically, $p(\mathbf{x}_{T})$ is a normal distribution $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The transformation from $p(\mathbf{x}_{T})$ to $p(\mathbf{x}_{0})$, usually known as reverse process, is an iterative process described by the learnable distribution $p_{\theta}(\mathbf{x}_{0:T})$:
$p_{\theta}(\mathbf{x}_{0:T}):=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),$ (1)
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}):=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$ (2)
where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ are latent variables with the same dimensionality as the samples of the target distribution that represent intermediate steps, and $\theta$ are the learnable parameters. This process can be seen as a Markovian chain, where a network learns Gaussian transitions $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ that predict the next latent variable $\mathbf{x}_{t-1}$ given $\mathbf{x}_{t}$ as input.
To construct a trainable objective for $\theta$, we also need to introduce the forward process, where noise is gradually added to a sample from our target distribution until it is indistinguishable from pure noise. It is defined as follows:
$\begin{split}q(\mathbf{x}_{1:T}|\mathbf{x}_{0})&:=\sum_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})\\ q(\mathbf{x}_{t}|\mathbf{x}_{t-1})&:=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})\end{split}$ (3)
where $\beta_{1},\dots,\beta_{T}$ are coefficients that schedule the amount of noise to be added at each step. An interesting property of the forward process is that we can express any arbitrary step $\mathbf{x}_{t}$ using the coefficients $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (4)
meaning that we can actually compute $\mathbf{x}_{t}$ without computing the entire chain:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\hskip 11.38092pt\text{where}\hskip 11.38092pt\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (5)
With the above forward procedure, Ho et al. Ho, Jain, and Abbeel (2020) proposed a simple loss function where the network directly learns the noise $\epsilon$ in the forward procedure:
$L_{\text{simple}}:=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$ (6)
Using a fixed variance for the reverse process to $\Sigma(\mathbf{x}_{t},t)~=~\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}}\beta_{t}\mathbf{I}$, the reverse step becomes:
$\mathbf{x}_{t-1}=\tilde{\mathbf{x}}_{0}^{(t)}\sqrt{\bar{\alpha}_{t-1}}+\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)\sqrt{1-\bar{\alpha}_{t-1}}$ (7)
where $\tilde{\mathbf{x}}_{0}^{(t)}$ is the denoised observation, a prediction of the generated sample $\mathbf{x}_{0}$ at the step $t$, which can be obtained by rearranging equation (5):
$\tilde{\mathbf{x}}_{0}^{(t)}:=\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)}{\sqrt{\bar{\alpha}_{t}}}$ (8)
II.2 Masked Sampling of Diffusion Models
The DDPM algorithm enables training a network $\epsilon_{\theta}$ that transforms Gaussian noise $\mathbf{x}_{T}$ to a high-fidelity flow field $\mathbf{x}_{0}$. In this section, we present a novel sampling method where the reverse process is further guided to generate a certain flow field corresponding to a given sparse observation.
The idea behind our method is based on the key observation that the sparse inputs $\mathbf{x_{s}}$ are authentic data points that should be preserved in the final generated sample. Thus, we introduce the sparse information $\mathbf{x_{s}}$ into the denoised observation $\tilde{\mathbf{x}}_{0}^{(t)}$ by applying a masking operation. Specifically, we define a binary mask $m\in\mathbb{R}^{n\times n}$, where entries corresponding to known ground truth values are set to 1, while all other positions are assigned a value of 0. After computing the denoised observation $\tilde{\mathbf{x}}^{(t)}_{0}$, we include the available true values via:
$\tilde{\mathbf{x}}^{(t)}_{m}:=\tilde{\mathbf{x}}^{(t)}_{0}\cdot(1-m_{t})+\mathbf{x_{s}}\cdot m_{t}$ (9)
where all the operations are performed element-wise. Since the prediction of $\mathbf{x}_{0}$ at the start of the reverse process is usually very poor, we introduce a scheduler $m_{t}~=~m~\cdot~\gamma_{t}$ where $\gamma_{t}\in[0,1]$ to regulate the strength of the mask. $\gamma_{t}$ is stronger when the prediction is nearly random (t is close to T) and weaker when we are closer to the solution (t is close to 0).
After masking the prediction with the real values, we can compute the next latent variable $\mathbf{x}_{t-1}$ following equation (7), substituting the denoised observation by our masked version $\tilde{\mathbf{x}}_{m}^{(t)}$:
$\mathbf{x}_{t-1}=\tilde{\mathbf{x}}_{m}^{(t)}\sqrt{\bar{\alpha}_{t-1}}+\epsilon_{\theta}(\mathbf{x}_{t},t)\sqrt{1-\bar{\alpha}_{t-1}}$ (10)
Starting the reverse process from pure noise $\mathbf{x}_{T}$, equation (10) will guide the model towards predicting a high-fidelity flow field that corresponds to the sparse observation $\mathbf{x}_{s}$.
Figure 1: a) Sparse data $\mathbf{x}_{s}$ marked with white crosses over ground truth. b) Smooth Gaussian mask $m^{s}$.
However, as described so far, the mask only yields single-point values onto the field, which can introduce discontinuities. Thus, we also propose a smooth Gaussian mask $m^{s}$. Given the set of sparse data points $P~=~\{\mathbf{p}^{1},\dots,\mathbf{p}^{K}\}$ representing the input $\mathbf{x_{s}}$, we define a function $f_{k}(\mathbf{x})$ for each point $\mathbf{p}^{k}$ that returns the value of a Gaussian centred at that point with a smoothing factor $\sigma$:
[…]
We define the Gaussian operator $m^{s}=G_{\mathbf{s}}(\mathbf{x_{s}})$, which computes this smooth mask $m^{s}$ by evaluating each function $f_{k}(\mathbf{x})$ across all cells of the mask. An example of a resulting mask can be seen in Figure 1(b). Meanwhile, since we have no true information of the values surrounding each mask point $\mathbf{p}^{k}$, we compute the interpolation between the true points using an interpolation operator $\mathbf{\hat{x}_{s}}=U_{\mathbf{s}}(\mathbf{x_{s}})$. Now, we can also mask the prediction $\tilde{\mathbf{x}}^{(t)}_{0}$ with intermediate values between the true sparse data following:
$\tilde{\mathbf{x}}^{(t)}_{m}=\tilde{\mathbf{x}}^{(t)}_{0}\cdot(1-m_{t}^{s})+\hat{\mathbf{x}}_{\mathbf{s}}\cdot m_{t}^{s}$ (12)
Similarly to the binary mask, we also apply a scheduler $m_{t}^{s}=m^{s}\gamma_{t}$ to the smooth mask. A comparison between the point-wise and the Gaussian masking methods can be found in Appendix A, which clearly displays the benefits of using the Gaussian mask. Thus, we continue with this masking method and we call it masked diffusion. Figure 2 illustrates the method, and a detailed summary of the procedure can be found in Algorithm 1.
Figure 2: Sketch of our proposed masked diffusion method, where a smooth mask is applied to improve the prediction of $\mathbf{x}_{0}$ during the reverse process.
Input: Sparse data sample $\mathbf{x_{s}}$
$m^{s}=G_{s}(\mathbf{x_{s}})$ ;
$\hat{\mathbf{x}}_{\mathbf{s}}=U_{s}(\mathbf{x}_{\mathbf{s}})$ ;
$\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ ;
for $t=T,\dots,1$ do
   $m^{s}_{t}=m^{s}\cdot\gamma_{t}$
  $\tilde{\mathbf{x}}_{0}^{(t)}=\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon_{\theta}(\mathbf{x}_{t},t)\right)/\sqrt{\bar{\alpha}_{t}}$ ;
  $\tilde{\mathbf{x}}^{(t)}_{m}=\tilde{\mathbf{x}}^{(t)}_{0}\cdot(1-m_{t}^{s})+\hat{\mathbf{x}}_{\mathbf{s}}\cdot m_{t}^{s}$ ;
  $\mathbf{x}_{t-1}=\tilde{\mathbf{x}}_{m}^{(t)}\sqrt{\bar{\alpha}_{t-1}}+\epsilon_{\theta}(\mathbf{x}_{t},t)\sqrt{1-\bar{\alpha}_{t-1}}$ ;
   end for
  return $\mathbf{x}_{0}$ ;
Algorithm 1 Masked diffusion algorithm
II.3 Physics-based Diffusion Training
The high fidelity samples $\mathbf{x}_{0}$ usually satisfy physical constraints, i.e., $\mathcal{F}(\mathbf{x})=0$, where $\mathcal{F}$ could be an underlying PDE stemming from conservation laws or representing boundary conditions. To enhance the physical coherence of the samples generated by DDPMs, we can integrate an additional physical residual loss into the training as:
$L_{\text{phy}}:=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}+\lambda\|\mathcal{F}(\tilde{\mathbf{x}}_{0}^{(t)})\|_{2}\right]$ (13)
where $\lambda$ is a balance factor for the physical residual. Similar methods have been proposed by previous works Bastek, Sun, and Kochmann (2024); Shan et al. (2024), such as the Physics-Informed Diffusion Models (PIDMs) Bastek, Sun, and Kochmann (2024). In PIDM, a central difficulty is that the scaling factor is highly dependent on the specificity of the training task at hand. We consider an alternative where the rescaling factor is dynamically set to balance the importance of each term of the global loss. This balance is computed via the ratio of the noise step and the physics residual loss as:
$\lambda_{\text{dyn}}=\frac{\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}}{\|\mathcal{F}(\tilde{\mathbf{x}}_{0}^{(t)})\|_{2}}$ (14)
To improve on these manually-scaled and heuristic approaches for balancing, we build on a recent method for conflict-free updates, ConFIG Liu, Chu, and Thuerey (2024), to automatically balance the loss terms. Defining $\boldsymbol{g}_{D}={\partial\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|_{2}}/{\partial\theta}$ and $\boldsymbol{g}_{F}={\partial\|\mathcal{F}(\tilde{\mathbf{x}}_{0}^{(t)})\|_{2}}/{\partial\theta}$ as the gradients for diffusion loss and physical residual loss terms, the ConFIG method returns a conflict-free update direction $\boldsymbol{g}_{c}$ as
$\begin{split}&\boldsymbol{g}_{v}=\mathcal{U}\left[\mathcal{U}\left(\mathcal{O}\left(\boldsymbol{g}_{D},\boldsymbol{g}_{F}\right)\right)+\mathcal{U}\left(\mathcal{O}\left(\boldsymbol{g}_{F},\boldsymbol{g}_{D}\right)\right)\right]\\ &\boldsymbol{g}_{c}=\left(\boldsymbol{g}_{D}^{\top}\boldsymbol{g}_{v}+\boldsymbol{g}_{F}^{\top}\boldsymbol{g}_{v}\right)\boldsymbol{g}_{v}\end{split}$ (15)
where $\mathcal{U}(\boldsymbol{g}_{i})=\boldsymbol{g}_{i}/|\boldsymbol{g}_{i}|+\varepsilon$ is the normalizing operator and $\mathcal{O}(\boldsymbol{g}_{i},\boldsymbol{g}_{j})=\boldsymbol{g}_{j}-\frac{\boldsymbol{g}_{i}^{\top}\boldsymbol{g}_{j}}{|\boldsymbol{g}_{i}|^{2}}\boldsymbol{g}_{i}$ the orthogonality operator.
Provably, $\boldsymbol{g}_{D}^{\top}\boldsymbol{g}_{c}>0$ and $\boldsymbol{g}_{F}^{\top}\boldsymbol{g}_{c}>0$ hold, meaning that optimization along $\boldsymbol{g}_{c}$ will decrease both the diffusion loss and physical residual loss simultaneously. The generalization of the operator to more than two gradients can be found in Liu’s et al. articleLiu, Chu, and Thuerey (2024). As shown in equation (15), the magnitude of $\boldsymbol{g}_{v}$ in the ConFIG method is rescaled based on the sum of each loss-specific gradient’s projection on the conflict-free direction. However, for sparse reconstruction tasks, the physical residual loss is usually significantly larger than the diffusion loss. This leads to the magnitude change of the $\boldsymbol{g}_{v}$ mainly tracking the magnitude change of the residual gradient. For momentum-based optimizers like Adam, this represents a problem, as these optimizers adaptively change the learning rate according to the variation of the gradient. In turn, it makes the balancing approach of the ConFIG method ineffective. To counteract this undesirable behavior, we propose the following modification to the ConFIG method, which computes the magnitude of the resulting gradient after normalizing both loss-specific gradients:
$\boldsymbol{g}_{c}=\left(\mathcal{U}(\boldsymbol{g}_{D})^{\top}\boldsymbol{g}_{v}+\mathcal{U}(\boldsymbol{g}_{F})^{\top}\boldsymbol{g}_{v}\right)\boldsymbol{g}_{v}$ (16)
This results in a method that considers both loss terms equally for the computation of the gradient magnitude. At the same time, it also allows the Adam optimizer to compute the moments of the gradient in a balanced manner. The caveat of this modification is that we do not transfer any information about the original magnitude of $\boldsymbol{g}_{D}$ and $\boldsymbol{g}_{F}$ to the optimizer, as the new magnitude only contains the information of how similar the direction of $\boldsymbol{g}_{v}$ is to the directions of both loss terms. Nonetheless, our results show that the benefits of conflict-free learning updates clearly outweigh the lack of information about the original magnitudes.
We call the resulting method $\text{ConFIG}_{u}$ in the following, and Figure 3 illustrates how our update strategy differs from the original ConFIG method for a specific two-dimensional example. The $\text{ConFIG}_{\text{u}}$ updates in blue preserve significantly more information from the diffusion updates and are generally more balanced. They effectively capture the agreement between the directions of the two loss gradients. We compare the performance of these training strategies in Section III.2.
(a) Conflicting gradients.
(b) Similarly directed gradients.
Figure 3: A comparison of ConFIG and $\text{ConFIG}_{\text{u}}$ methods, where $\boldsymbol{g}_{F}$ is the gradient of the physical loss and $\boldsymbol{g}_{D}$ the gradient of the diffusion loss.
III Results and discussion
In this section, we evaluate our method in two challenging scenarios: 2D Kolmogorov flow and 3D isotropic turbulence. We first compare the performance of different training methods, demonstrating how our $\text{ConFIG}_{\text{u}}$ training method outperforms existing approaches in generating physically consistent samples. Next, we investigate the effectiveness of our proposed masked diffusion sampling method for sparse data reconstruction. The results show that our method effectively reconstructs fine-grained details and preserves the structural integrity of the flow fields, outperforming baseline methods, especially in capturing high-frequency components that are critical in turbulent flows. Finally, we provide a qualitative and quantitative analysis of the reconstructed samples, including visualizations and an evaluation of the vorticity and Q-criterion distributions.
III.1 Test cases
(a) Examples of the 2D dataset, comprised of vorticity field snapshots of DNS simulations.
(b) Example of one snapshot of the 3D dataset. We plot 2D z-slices through the middle of the domain for the three velocity components.
Figure 4: Datasets used: (a) 2D Kolmogorov dataset, (b) 3D isotropic turbulence.
For the 2D Kolmogorov flow, we follow the configuration of the previous work Shu, Li, and Farimani (2023). The system is governed by the following Navier-Stokes equation in its vorticity form:
$\frac{\partial\omega(\mathbf{x},t)}{\partial t}+\mathbf{u}(\mathbf{x},t)\cdot\nabla\omega(\mathbf{x},t)=\frac{1}{Re}\nabla^{2}\omega(\mathbf{x},t)+f(\mathbf{x})$ (17)
where $\textbf{x}\in(0,2\pi)^{2}$, $\omega$ is the vorticity, $\mathbf{u}$ the velocity field and $f(\mathbf{x})=-4\cos(4x_{2})-0.1\omega(\mathbf{x},t)$ is the external force. The Reynolds number $Re$ is set to 1000. The simulation data contains 4240 samples of size $256\times 256$, from which 3000 are randomly selected for training and the remaining 1240 to test the models. To evaluate the physical consistency of the generated samples, we compute the residual of equation (17), where the time derivative is approximated with three generated consecutive timesteps as $\partial_{t}~\omega~\approx~(\omega(\mathbf{x},t-\Delta t)+\omega(\mathbf{x},t+\Delta t))/(2\Delta t)$, and the spatial derivates are calculated with the Fourier pseudo-spectral method.
For the 3D turbulent scenario, we choose a forced isotropic turbulence simulation from the Johns Hopkins Turbulence Database (JHTDB) Databases (2019). It consists of a DNS simulation with a pseudo-spectral method using 10243 nodes in a periodic domain of size $(0,2\pi)^{3}$ during $T=10$ s using 5028 timesteps. Energy is injected to keep the total energy constant in modes such that their wave number magnitude is less than or equal to 2. We use a smaller derived dataset, taking 500 snapshots at a smaller resolution of $128^{3}$ using spectral downsampling. Finally, we shuffle all snapshots and use 400 and 100 samples for training and testing, respectively. The divergence of the flow field $\nabla\cdot\mathbf{u}$ is used as the metric for physics consistency. Figures 4(a) and 4(b) show samples of our training data.
III.2 Physcial-consistency
We first target an unconditioned generation task to examine the ability of different methods to generate physically constrained samples. We consider standard diffusion models, Physics-Informed Diffusion Models Bastek, Sun, and Kochmann (2024) with dynamic rescale factor ($\text{PIDM}_{\text{dyn}}$), ConFIG Liu, Chu, and Thuerey (2024), and our $\text{ConFIG}_{\text{u}}$. Some comments about training hyperparameters and the network’s architecture can be found in Appendix B.
To calculate the physical residual, the time and spatial derivatives are evaluated in a discrete manner. Ideally, the training data’s physical residual $\|\mathcal{R}(\omega^{*})\|_{2}$ should be 0. Still, we usually encounter a non-zero physical residual due to the discretisation errors. As no method can be expected to completely resolve this discretization error, it obscures the capabilities of the methods to improve the diffusion training. Hence, we employ a metric that compares the residual of the generated samples $\omega^{\prime}$ to the average of the test dataset samples:
$\boldsymbol{L_{\text{{res}}}}=|\|\mathcal{R}(\omega^{\prime})\|_{2}-\mathcal{R}_{\text{avg}}^{*}|$ (18)
where $\mathcal{R}_{\text{avg}}^{*}:=\frac{1}{N}\sum_{i}^{N}\|\mathcal{R}(\omega^{*})\|_{2}$ is the average physical residual of the test dataset.
Test case Training method $\boldsymbol{L_{\text{{res}}}}$ $\|\mu^{\prime}-\mu^{*}\|_{2}$
2D Kolmogorov flow Std. Diffusion 1.27 $\pm$ 0.12 0.455 $\pm$ 0.104
$\text{PIDM}_{\text{dyn}}$ Bastek, Sun, and Kochmann (2024) 0.99 $\pm$ 0.21 0.316 $\pm$ 0.005
ConFIG Liu, Chu, and Thuerey (2024) 1.24 $\pm$ 0.32 0.340 $\pm$ 0.009
$\text{ConFIG}_{\text{u}}$ 0.68 $\pm$ 0.15 0.309 $\pm$ 0.008
3D isotropic turbulence Std. Diffusion 1.03 $\pm$ 0.07 0.690 $\pm$ 0.037
$\text{PIDM}_{\text{dyn}}$ Bastek, Sun, and Kochmann (2024) 1.08 $\pm$ 0.07 0.696 $\pm$ 0.042
ConFIG Liu, Chu, and Thuerey (2024) 1.01 $\pm$ 0.05 0.667 $\pm$ 0.067
$\text{ConFIG}_{\text{u}}$ 0.80 $\pm$ 0.05 0.663 $\pm$ 0.046
Table 1: Evaluation of the training methods under consideration. $\mu^{\prime}$ and $\mu^{*}$ represent the pixel-wise mean of the generated and the test samples, respectively.
[…]
III.3 Reconstruction tasks
Next, we test the proposed masked diffusion method for reconstructing high-fidelity flow fields from sparse data points. Following the configuration from the previous study by Shu et al. Shu, Li, and Farimani (2023), we target reconstruction tasks with sparse observations of 5% and 1.5625% of the high fidelity data.
The reconstructed samples should meet three key properties: accurate point-wise values, physical coherence, and a similar fluid structure to the ground truth.
The first property is typically measured using the Root Mean Square Error (RMSE) between the predicted flow field and the ground truth. To satisfy physical coherence, the reconstructed field should follow the physical residual of the data, which can be evaluated using the relative residual loss $\boldsymbol{L_{\text{{res}}}}$ introduced in Section III.2.
Evaluating the structure of a fluid flow improves our understanding of how closely the predicted sample matches the target high-fidelity field. As no ground truth is available in this probabilistic setting, we use a learned, perceptual similarity metric (LSiM) Kohl, Um, and Thuerey (2020) to measure the similarity between our reconstruction and the ground truth. This metric allows us to assess whether the structural characteristics of the generated sample align with those of the target field.
We also use the error in the energy spectrum to analyse errors in the frequency domain. For the 3D case, we compute the L1 integral norm in logarithmic scale of the difference between the ground truth and the reconstructed spectra, given by:
$\int_{k_{min}}^{k_{max}}\left|\log E(k)-\log E_{GT}(k)\right|d(\log k)$ (19)
For the 2D scenario, we compute this error using the enstrophy spectrum instead of the energy spectrum as used in other works Chen et al. (2024), since enstrophy cascades to small scales in 2D and therefore provides a more sensitive and physically meaningful diagnostic of whether the simulation captures the correct multiscale turbulent dynamics.
The smoothing coefficient of the mask $\sigma$ is a hyperparameter in our method. It determines the width of the Gaussian functions of the mask, centred at every value of the sparse data. We conducted a parameter exploration, shown in Figure 5 and Figure 6, which allowed us to obtain the values of $\sigma=0.038$ and $\sigma=0.052$ for the 5% and 1.5625% tasks in 2D, respectively, and $\sigma=0.06$ and $\sigma=0.08$ in 3D. Meanwhile, we additionally use a scheduler to control the strength of the mask in our sampling approach, as discussed in Section II.2. Experimentally, we observed that the scheduler $\gamma_{t}=\left(t/T\right)^{3}$ gave satisfactory results.
(a) 1.5625% task, sigma fixed at $\sigma=0.052$.
(b) 5% task, sigma fixed at $\sigma=0.038$.
Figure 5: Sigma value exploration for the 2D Kolmogorov case. The red vertical line indicates the chosen value for the parameter. All metrics are normalised.
(a) 1.5625% task, sigma fixed at $\sigma=0.08$
(b) 5% task, sigma fixed at $\sigma=0.06$
Figure 6: Sigma value exploration for the 3D isotropic turbulence case. All metrics are normalised.
Besides comparing with the method of Shu et al. Shu, Li, and Farimani (2023), we also compare our reconstruction result with the supervised baseline where the network directly predicts the high-fidelity data from an interpolated sparse input. For a fair comparison, we use the same backbone architectures as the ones for the diffusion models. Moreover, we need to highlight that the baseline and other diffusion models for sparse data reconstruction require specific training for each new scenario (different % of data, for instance) to maximize their performance. In contrast, the masked diffusion method is general and only needs to be trained once, as an unconditional generator, using only the high-resolution training samples. Table 2 shows the results of the 2D 5% reconstruction task, where we can observe how our masked diffusion method outperforms previous methods on the point-wise, residual, structure similarity, and energy spectrum errors. Similar results can be observed when evaluating the 1.5625% task, which can be seen in Table 3. Our method consistently obtains the lowest errors in both tasks.
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.79 $\pm$ 0.09 0.16 $\pm$ 0.07 0.29 $\pm$ 0.02 0.46 $\pm$ 0.12
Shu et al. Shu, Li, and Farimani (2023) 0.18 $\pm$ 0.03 0.84 $\pm$ 0.07 0.10 $\pm$ 0.02 0.54 $\pm$ 0.08
Masked diffusion (ours) 0.13 $\pm$ 0.02 0.02 $\pm$ 0.03 0.05 $\pm$ 0.01 0.26 $\pm$ 0.05
Table 2: Evaluation of the reconstruction task from 5% of the data for 2D Kolmogorov flow.
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 1.06 $\pm$ 0.16 0.21 $\pm$ 0.14 0.42 $\pm$ 0.06 0.68 $\pm$ 0.23
Shu et al. Shu, Li, and Farimani (2023) 0.37 $\pm$ 0.06 0.22 $\pm$ 0.11 0.17 $\pm$ 0.03 0.52 $\pm$ 0.12
Masked diffusion (ours) 0.31 $\pm$ 0.06 0.15 $\pm$ 0.06 0.11 $\pm$ 0.02 0.33 $\pm$ 0.07
Table 3: Evaluation of the reconstruction task from 1.5625% of the data for 2D Kolmogorov flow.
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.13 $\pm$ 0.01 0.32 $\pm$ 0.02 0.05 $\pm$ 0.01 0.64 $\pm$ 0.02
Shu et al. Shu, Li, and Farimani (2023) 0.45 $\pm$ 0.03 0.80 $\pm$ 0.06 0.64 $\pm$ 0.15 0.38 $\pm$ 0.07
Masked diffusion (ours) 0.17 $\pm$ 0.01 0.61 $\pm$ 0.03 0.06 $\pm$ 0.01 0.09 $\pm$ 0.01
Table 4: Evaluation of the reconstruction task from 5% of the data for the 3D turbulent case.
Models RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Supervised baseline 0.16 $\pm$ 0.01 0.36 $\pm$ 0.01 0.08 $\pm$ 0.01 0.71 $\pm$ 0.02
Shu et al. Shu, Li, and Farimani (2023) 0.44 $\pm$ 0.02 0.79 $\pm$ 0.05 0.62 $\pm$ 0.11 0.36 $\pm$ 0.04
Masked diffusion (ours) 0.20 $\pm$ 0.01 0.66 $\pm$ 0.03 0.09 $\pm$ 0.01 0.11 $\pm$ 0.01
Table 5: Evaluation of the reconstruction task from 1.5625% of the data for the 3D turbulent case.
In Tables 4 and 5 we show the results for the 3D case of the same two reconstruction tasks. Our masked diffusion method outperforms Shu et al.’s method across all evaluation metrics. However, the supervised baseline achieves lower errors in RMSE, physical residual, and LSiM. These three metrics are primarily influenced by the accuracy of the low-frequency components, which the supervised baseline captures more effectively. As shown in Figure 7, only the diffusion methods are capable of accurately reconstructing the high-frequency components. In contrast, the supervised baseline produces noticeably blurrier results despite achieving a lower RMSE. The spectrum in Figure 7 shows that the high wavenumber content of the solutions inferred by the supervised baseline clearly lacks energy.
(a) 5% sparse data reconstruction.
(b) 1.5625% sparse data reconstruction.
Figure 7: Energy spectrum comparison of different reconstruction methods for our two 3D scenarios.
Figures 8-11 show samples of the reconstruction in 2D and 3D cases. It can be seen how masked diffusion better approximates the complicated structures of the turbulent flow fields, producing sharper images and yielding a closer approximation of the flow structures and statistics.
Figure 8: Example 2D Kolmogorov reconstruction 5%
Figure 9: Example 2D Kolmogorov reconstruction 1.5625%
Figure 10: Example 3D isotropic turbulence 5% sparse data reconstruction (2D slices).
Figure 11: Example 3D isotropic turbulence 1.5625% sparse data reconstruction (2D slices).
The advantage of the supervised baseline stems from its training strategy: the network learns to interpolate missing data points directly from sparse observations. This interpolation performs well for large-scale structures but fails to reliably capture small-scale features. However, reconstructing high-frequency components is particularly important in turbulent simulations, where dynamics span a wide range of physical scales. For this reason, we consider the spectrum error as the central performance metric.
As is evident from the spectrum errors in Tables 4 and 5, only the two considered diffusion models are capable of capturing small-scale features. E.g., our method features more than $6.4\times$ lower error than the supervised baseline. However, the two diffusion approaches behave very differently when considering different amounts of available data. In Figure 12, we plot the RMSE over the 3D reconstruction of 50 samples with Shu et al.’s method and our masked diffusion method together with cubic interpolation. We observe that Shu et al.’s method remains mostly constant, with a consistently higher RMSE than our method. We see that if the percentage of available data is too low, our method achieves even lower RMSE errors than cubic interpolation, in addition to producing high-frequency content.
Figure 12: RMSE of the 3D isotropic flow reconstruction with varying percentages of sparse data.
We also study the distributional accuracy of the methods. For the 2D experiment, we measure the distribution of the vorticity magnitude of the reconstructed samples, as shown in Figure 13. Our method closely approximates the reference’s vorticity distribution on the 5% task, while Shu et al.’s method slightly underestimates the magnitude of the generated samples Shu, Li, and Farimani (2023). When evaluating the 1.5625% task, our method also underestimates the vorticity, but gives a better distribution than Shu et al.’s method.
For the 3D case, we compare the probability distribution of the Q-criterion, which is a scalar field used in fluid mechanics to identify vortices in a flow field. The definition of the $Q$-criterion is based on the decomposition of the velocity gradient into a symmetric part (strain rate tensor $S$) and an antisymmetric part (rotation tensor $\Omega$), and is computed as:
$Q=\tfrac{1}{2}\left(\|\Omega\|^{2}-\|S\|^{2}\right)$ (20)
$S_{ij}=\tfrac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right),\hskip 5.69046pt\Omega_{ij}=\tfrac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{j}}{\partial x_{i}}\right)$ (21)
In Figure 14, we clearly observe that our masked diffusion is the closest to the ground truth, while the supervised baseline overestimates the zero component. Figures 15 and 16 show visualizations of the Q-criterion, highlighting how our proposed method is the only one that captures most of the smaller vortices in the fluid.
When discussing efficiency, our method obtains these results using fewer reverse steps than Shu et al.Shu, Li, and Farimani (2023), which makes it a more computationally efficient model. For instance, for the more demanding 3D experiments in a NVIDIA L40S GPU, as shown in Appendix C, 100 reverse steps, i.e. NN evaluations, are enough to achieve maximum performance, compared to the T=1000 during training, or the 300 reverse steps used with Shu et al.’s method. This also results in shorter running times, where the supervised baseline needs around 0.04 s for a single sample, Shu et al.’s method 25 s, and our method 7.7 s. The only inconvenience of our method would be higher memory demands. During masked diffusion, we need to store the mask and the interpolated sparse data, resulting in 5.2 GB, compared with the 4.0 GB of Shu et al.’s method, and 1.4 GB of the single-step supervised baseline.
Overall, our proposed training and sampling techniques demonstrate a clearly improved performance compared to existing approaches. These results highlight the ability of our masked diffusion framework to capture fine-grained details and preserve the structural integrity of the reconstructed fluid fields.
(a) 5% sparse data reconstruction.
(b) 1.5625% sparse data reconstruction.
Figure 13: 2D Vorticity magnitude distribution of reconstructed samples.
(a) 5% sparse data reconstruction.
(b) 1.5625% sparse data reconstruction.
Figure 14: Q-criterion distribution of reconstructed samples.
Figure 15: Q-citerion in 3D for the 5% sparse data reconstruction. 3D plots showing Q isosurfaces with vorticity coloring.
Figure 16: Q-citerion in 3D for the 1.5625% sparse data reconstruction. 3D plots showing Q isosurfaces with vorticity coloring.
IV Conclusions
In this work, we proposed a novel method for reconstructing high-fidelity flow fields from limited sparse data using diffusion models. Instead of generating a random sample of the target distribution, our technique utilises the available sparse information to guide the sampling process towards our desired reconstructed sample. This results in trained models that better approximate the fluid’s structure, as they generate the missing information utilising the available data as a reference for the overall structure. Moreover, we introduced the $\text{ConFIG}_{\text{u}}$ training method to ensure physically accurate solutions, and showed how multiple targets can be combined in a non-destructive manner.
We have shown that our method outperforms previous diffusion-based reconstruction approaches in both 2D and 3D scenarios. In particular, it excels at producing physically consistent reconstructions while preserving key flow statistics and accurately capturing the smallest scales.
Our results point to several interesting avenues for future work, such as evaluating how the proposed generative method performs for reconstructing other flow problems. Additionally, the generalization capabilities of our trained models to different Reynolds numbers and boundary conditions, as well as robustness to noise, are a promising avenues for future studies.
We believe that our efforts demonstrate the promise of diffusion models for highly accurate reconstructions of physical flow fields given sparse measurements, an important open problem in different fields of engineering. Our work also highlights the importance of non-conflicting algorithms when training generative models with multiple objectives and physics constraints.
Our approach generally enourages further research into physics-constrained diffusion models for addressing probabilistic flow-related problems.
Acknowledgements.
This work was partially supported by the ERC Consolidator Grant SpaTe (CoG-2019-863850) and by the China Scholarship Council (No.202206120036 ).
Data Availability Statement
The data and source code that support the findings of this study are openly available in GitHub at https://github.com/tum-pbs/sparse-reconstruction.
Author Declarations
Author Declarations
The authors have no conflicts to disclouse.
Author contributions
Marc Amorós-Trepat: Conceptualization (equal); Methodology (equal); Investigation (equal); Data Curation (equal); Formal Analysis (equal); Software (lead); Validation (equal); Visualization (equal); Writing – Original Draft (equal); Writing – Review & Editing (equal); Luis Medrano-Navarro: Investigation (equal); Data Curation (equal); Formal Analysis (equal); Software (equal); Validation (equal); Visualization (equal); Writing – Original Draft (equal); Writing – Review & Editing (equal); Qiang Liu: Conceptualization (equal); Methodology (equal); Writing – Original Draft (equal); Writing – Review & Editing (equal); Supervision (equal); Luca Guastoni: Writing – Review & Editing (equal); Supervision (equal); Nils Thuerey: Writing – Review & Editing (equal); Supervision (equal);
Appendix A Comparison between point-wise and Gaussian masking methods
In Section II.2, we introduced a masking strategy to improve the prediction of the resulting high-fidelity field at any given timestep $t$. We proposed two variants of this mask: one that activates only at specific points where the ground truth value is known, and another that incorporates a Gaussian smoothing function around each known value. Table 6 compares these methods with the task of reconstructing a flow field from 1.5625% of the data, showing how the Gaussian mask consistently outperforms the point-wise mask across all metrics introduced in the Results section.
Methods RMSE $\|\mathcal{R}(\tilde{\omega})\|_{2}$ $\boldsymbol{L_{\text{{res}}}}$ LSiM
Interpolation 0.4930 106.2809 102.9909 0.1552
Point-wise Mask 0.3500 4.2134 0.9234 0.1645
Gaussian Mask 0.3024 3.5658 0.2758 0.1094
Table 6: Comparison between nearest interpolation, and our proposed masking methods. The evaluated task is the reconstruction of a field from 1.5625% of the known values. The used model was trained with the $\text{ConFIG}_{u}$ method.
Appendix B Training details
To obtain comparable results, all 2D models were trained for 1000 epochs and a batch size of 5. A learning rate of $10^{-4}$ was used, with an exponential scheduler with $\gamma=0.98$, and a minimum $lr=10^{-5}$. The training of the 3D models was slightly different. Due to having a smaller number of samples we trained during 2000 epochs to achieve convergence. Since the samples are much bigger we needed a bigger GPU (NVIDIA L40S with 50 GB) for training and a slightly different batch size (4). We still used the same variable learning rate. The noise schedule $\beta$ for the diffusion forward process was kept as linear between $10^{-4}$ and 0.02.
The 2D network is based on a UNet model Ronneberger, Fischer, and Brox (2015), which is a convolutional neural network architecture that features a symmetric encoder–decoder structure with skip connections to capture both global context and fine details. To further expand the receptive field of the network, we also use self-attentionVaswani (2017) in the bottleneck layer.
As backbone model for 3D fields, we use the P3D model Holzschuh et al. (2025), a recent hybrid CNN-Transformer architecture for 3D physical simulations that can be applied in patches or to the whole sample. It follows a typical hierarchical U-shaped structure with an encoder and a decoder, based on a combination of convolutional and transformer blocks, which again combine the ideas from Swin transformers Liang et al. (2021) and diffusion transformers (DiT) Peebles and Xie (2023).
Appendix C Ablation of DDIM reverse steps
We optimized the choice of the number of reverse steps in DDIM. In Figure 17, for the 3D case, the metrics stabilize after around 100 steps. A similar result is obtained in the 2D case. Thus, we chose this value for the final experiments.
Figure 17: Optimization of DDIM reverse steps. Evolution of three normalized metrics for an increasing number of reverse steps.